How much salt must be added to 60 Ltr. of a 20% solution of salt to increase it to 40% solution of salt?
Answer by richwmiller(17219)

x+.2*60=.4(x+60)
x+.2*60=.4x+.4*60
x+12=.4x+24
.6x=12
6x=120
x=20 liters of pure salt